# Global well-posedness for the generalized intermediate NLS with a nonvanishing condition at infinity

### v1

@@ 2.1. Notation
We have already defined the space $\mathcal{Z}^{k}_{\rho}$, the distance function $d^{k}_{\rho}(\cdot,\cdot)$ and the functional $\mathcal{E}^{k}_{\rho}(\cdot)$ on it, in (1.2), (1.3) and (1.4), respectively. We also define the smooth subset $\mathcal{Z}^{\infty}_{\rho}:=\bigcap_{k\geq 1}\mathcal{Z}^{k}_{\rho}$.
@@ 2.2. Properties of $T_{\delta}$
Having an unbounded symbol (1.1), the operator $T_{\delta}$ itself is not bounded on $L^{p}(\mathbb{R})$ nor on $H^{s}(\mathbb{R})$. However, since the Fourier symbol $r_{\delta}(\xi)=(\coth(\delta\xi)-\mathrm{sgn}(\xi))\xi$ of $\mathcal{L}_{\delta}$ satisfies $|r_{\delta}(\xi)|\lesssim\delta^{-1}$, by Plancherel’s theorem we see that $\mathcal{L}_{\delta}$ is bounded on $H^{s}(\mathbb{R})$ for any $s\in\mathbb{R}$ and satisfies
@@
For $u\in\mathcal{Z}^{k}_{\rho}$, we notice that $\partial_{x}|u|^{2}=\partial_{x}(|u|^{2}-\rho^{2})$ gives
(2.4) $\|T_{\delta}\partial_{x}|u|^{2}\|_{L^{2}}$ $\lesssim\delta^{-1}\big\||u|^{2}-\rho^{2}\big\|_{L^{2}}+\|u\|_{L^{\infty}}\|\partial_{x}u\|_{L^{2}},$
$\|\mathcal{L}_{\delta}|u|^{2}\|_{L^{2}}$ $\lesssim\delta^{-1}\big\||u|^{2}-\rho^{2}\big\|_{L^{2}}.$
@@
2.3. Properties of the space $\mathcal{Z}^{k}_{\rho}$ and control of Littlewood-Paley operators
We begin this subsection with a basic estimate in $\mathcal{Z}^{1}_{\rho}$, which will be used in Section 4:
(2.6) $\rho\leq\|f\|_{L^{\infty}}\lesssim\rho+\big\||f|^{2}-\rho^{2}\big\|_{L^{2}}^{\frac{1}{3}}\|\partial_{x}f\|_{L^{2}}^{\frac{1}{3}}\qquad(f\in\mathcal{Z}^{1}_{\rho}).$
The first inequality is clear and follows from the definition of $\rho$ and $\|\cdot\|_{L^{\infty}}$. The second inequality is obtained by Gagliardo-Nirenberg and Young’s inequalities as follows:
@@ 2.3. Properties of the space $\mathcal{Z}^{k}_{\rho}$ and control of Littlewood-Paley operators
Next, from Sobolev’s embedding $H^{1}(\mathbb{R})\hookrightarrow L^{\infty}(\mathbb{R})$ we see that for any integer $k\geq 1$, $f\in\mathcal{Z}^{k}_{\rho}$ and $g\in H^{k}(\mathbb{R})$, it holds $f+g\in\mathcal{Z}^{k}_{\rho}$ and
(2.7) $\mathcal{E}^{k}_{\rho}(f+g)\lesssim\mathcal{E}^{k}_{\rho}(f)+\|g\|_{H^{k}}^{2}.$
In a similar manner (but using the Gagliardo-Nirenberg inequality instead of Sobolev’s embedding), for $f_{1},f_{2}\in\mathcal{Z}^{1}_{\rho}$ and $g_{1},g_{2}\in H^{1}(\mathbb{R})$, we can show that
(2.8) $d^{1}_{\rho}(f_{1}+g_{1},f_{2}+g_{2})$ $\lesssim\big(1+\|g_{1}\|_{L^{2}}\big)d^{1}_{\rho}(f_{1},f_{2})$
@@
Below, we collect some estimates for frequency localized functions. They are all easily proven using Bernstein’s inequality. For any $\ell\geq 0$ and $k,j\geq 1$, we can show that $P_{\leq\ell}:\mathcal{Z}^{k}_{\rho}\to\mathcal{Z}^{k+j}_{\rho}$ with the estimates
(2.9) $\begin{aligned} \mathcal{E}^{k}_{\rho}(P_{\leq\ell}f)&\leq\mathcal{E}^{k}_{\rho}(f)+\big\||P_{\leq\ell}f|^{2}-|f|^{2}\big\|_{L^{2}}\leq\mathcal{E}^{k}_{\rho}(f)+2\|f\|_{L^{\infty}}\|P_{\geq\ell+1}f\|_{L^{2}}\\ &\leq\mathcal{E}^{k}_{\rho}(f)+2^{1-\ell}\|f\|_{L^{\infty}}\|\partial_{x}f\|_{L^{2}}\leq\mathcal{E}^{k}_{\rho}(f)+2^{1-\ell}\mathcal{E}^{1}_{\rho}(f)\leq 3\mathcal{E}^{k}_{\rho}(f),\end{aligned}$
(2.10) $\mathcal{E}^{k+j}_{\rho}(P_{\leq\ell}f)\lesssim 2^{j\ell}\mathcal{E}^{k}_{\rho}(f)+2^{-\ell}\mathcal{E}^{1}_{\rho}(f)\lesssim 2^{j\ell}\mathcal{E}^{k}_{\rho}(f).$
We also see that $P_{\leq\ell}f\to f$ in $\mathcal{Z}^{k}_{\rho}$ as $\ell\to\infty$, thanks to the estimate
(2.11) $d^{k}_{\rho}(P_{\leq\ell}f,f)$ $\leq\|P_{\geq\ell+1}\partial_{x}f\|_{H^{k-1}}+\|P_{\geq\ell+1}f\|_{L^{\infty}}+\big\||P_{\leq\ell}f|^{2}-|f|^{2}\big\|_{L^{2}}$
@@
Note that if $f\in C([0,T];\mathcal{Z}^{k}_{\rho})$, then the above convergence is uniform in $t$. Using the identity
(2.12) $|P_{\leq\ell}f|^{2}-|P_{\leq\ell}g|^{2}$
$=|f|^{2}-|g|^{2}-\Big(P_{\geq\ell+1}f\cdot\overline{P_{\leq\ell}f}+f\cdot\overline{P_{\geq\ell+1}f}\Big)+\Big(P_{\geq\ell+1}g\cdot\overline{P_{\leq\ell}g}+g\cdot\overline{P_{\geq\ell+1}g}\Big)$
$=|f|^{2}-|g|^{2}-P_{\geq\ell+1}(f-g)\cdot\overline{P_{\leq\ell}f}-P_{\geq\ell+1}g\cdot\overline{P_{\leq\ell}(f-g)}-(f-g)\overline{P_{\geq\ell+1}f}-g\overline{P_{\geq\ell+1}(f-g)},$
we have
(2.13) $\begin{aligned} d^{k}_{\rho}(P_{\leq\ell}f,P_{\leq\ell}g)&\leq d^{k}_{\rho}(f,g)\begin{aligned} &+\|P_{\geq\ell+1}(f-g)\|_{L^{2}}\|P_{\leq\ell}f\|_{L^{\infty}}+\|P_{\geq\ell+1}g\|_{L^{2}}\|P_{\leq\ell}(f-g)\|_{L^{\infty}}\\ &+\|f-g\|_{L^{\infty}}\|P_{\geq\ell+1}f\|_{L^{2}}+\|g\|_{L^{\infty}}\|P_{\geq\ell+1}(f-g)\|_{L^{2}}\end{aligned}\\ &\leq d^{k}_{\rho}(f,g)+2^{1-\ell}\Big(\big[\mathcal{E}^{1}_{\rho}(f)\big]^{\frac{1}{2}}+\big[\mathcal{E}^{1}_{\rho}(g)\big]^{\frac{1}{2}}\Big)d^{1}_{\rho}(f,g),\end{aligned}$
and
(2.14) $d^{k+j}_{\rho}(P_{\leq\ell}f,P_{\leq\ell}g)\lesssim 2^{j\ell}d^{k}_{\rho}(f,g)+2^{-\ell}\Big(\big[\mathcal{E}^{1}_{\rho}(f)\big]^{\frac{1}{2}}+\big[\mathcal{E}^{1}_{\rho}(g)\big]^{\frac{1}{2}}\Big)d^{1}_{\rho}(f,g).$
Finally, for different $\ell,m\in\mathbb{Z}_{\geq 0}$ and $f,g\in\mathcal{Z}^{2}_{\rho}$, we use (2.11) with $k=1$ to see that
(2.15) $d^{1}_{\rho}(P_{\leq\ell}f,P_{\leq m}g)\lesssim d^{1}_{\rho}(f,g)+2^{-\ell}\Big(\big[\mathcal{E}^{2}_{\rho}(f)\big]^{\frac{1}{2}}+\mathcal{E}^{2}_{\rho}(f)\Big)+2^{-m}\Big(\big[\mathcal{E}^{2}_{\rho}(g)\big]^{\frac{1}{2}}+\mathcal{E}^{2}_{\rho}(g)\Big).$
@@ 2.4. Estimates for the free Schrödinger operator
The action of $U(t)$ in the Zhidkov-type spaces has been well-studied; see, e.g., [Gerard, Chen].
@@
Using Gagliardo-Nirenberg inequality, we see that
@@
(2.19) $\phi\in\mathcal{Z}^{k}_{\rho}\quad\Longrightarrow\quad U(t)\phi\in C(\mathbb{R};\mathcal{Z}^{k}_{\rho}),\quad\mathcal{E}^{k}_{\rho}(U(t)\phi)\lesssim(1+|t|^{\frac{3}{4}})\mathcal{E}^{k}_{\rho}(\phi).$
@@
(2.20) $d^{k}_{\rho}(U(t)\phi,U(t)\psi)\lesssim(1+|t|^{\frac{3}{4}})\Big(1+\mathcal{E}^{1}_{\rho}(\phi)+\mathcal{E}^{1}_{\rho}(\psi)\Big)^{\frac{1}{2}}d^{k}_{\rho}(\phi,\psi)$
for any $k\geq 1$ and $\phi,\psi\in\mathcal{Z}^{k}_{\rho}$. To prove (2.20), it is clearly sufficient to estimate the non-Sobolev part of $d^{1}_{\rho}$. From (2.17), we have
$\|U(t)\phi-U(t)\psi\|_{L^{\infty}}\lesssim\|\phi-\psi\|_{L^{\infty}}+|t|^{\frac{1}{4}}\|\partial_{x}(\phi-\psi)\|_{L^{2}}\lesssim(1+|t|^{\frac{1}{4}})d^{1}_{\rho}(\phi,\psi).$
Using the decomposition similar to (2.12) and (2.16), (2.17), we have
$\big\||U(t)\phi|^{2}-|U(t)\psi|^{2}\big\|_{L^{2}}$
$\lesssim\big\||\phi|^{2}-|\psi|^{2}\big\|_{L^{2}}\begin{aligned} &+\big\|U(t)(\phi-\psi)-(\phi-\psi)\big\|_{L^{2}}\|U(t)\phi\|_{L^{\infty}}+\|U(t)\psi-\psi\|_{L^{2}}\|U(t)(\phi-\psi)\|_{L^{\infty}}\\ &+\|\phi-\psi\|_{L^{\infty}}\|U(t)\phi-\phi\|_{L^{2}}+\|\psi\|_{L^{\infty}}\big\|U(t)(\phi-\psi)-(\phi-\psi)\big\|_{L^{2}}\end{aligned}$
$\lesssim(1+|t|^{\frac{3}{4}})\Big(1+\mathcal{E}^{1}_{\rho}(\phi)+\mathcal{E}^{1}_{\rho}(\psi)\Big)^{\frac{1}{2}}d^{1}_{\rho}(\phi,\psi),$
@@ 1. $\mathrm{(i)}$
    For $f\in\mathcal{Z}^{1}_{\rho}$,
    $\|\mathcal{N}_{\delta}(f)\|_{L^{2}}\lesssim\delta^{-1}\big[\mathcal{E}^{1}_{\rho}(f)\big]^{\frac{3}{2}}.$
    For any integer $k\geq 1$ and $f\in\mathcal{Z}^{k}_{\rho}$,
    $\|\partial_{x}^{k}\mathcal{N}_{\delta}(f)\|_{L^{2}}\lesssim_{k}\delta^{-1}\mathcal{E}^{1}_{\rho}(f)\|\partial_{x}f\|_{H^{k-1}}.$
@@ 2. $\mathrm{(ii)}$
    For any integer $k\geq 0$ and $f\in\mathcal{Z}^{k+1}_{\rho}$,
    $\|\partial_{x}^{k}\mathcal{N}_{\infty}(f)\|_{L^{2}}\lesssim_{k}\mathcal{E}^{1}_{\rho}(f)\|\partial_{x}f\|_{H^{k}}.$
@@ 3. $\mathrm{(iii)}$
    For $f_{1},f_{2}\in\mathcal{Z}^{1}_{\rho}$,
    $\delta\;\|\mathcal{N}_{\delta}(f_{1})-\mathcal{N}_{\delta}(f_{2})\|_{H^{1}}+\|\mathcal{N}_{\infty}(f_{1})-\mathcal{N}_{\infty}(f_{2})\|_{L^{2}}\lesssim\Big(1+\mathcal{E}^{1}_{\rho}(f_{1})+\mathcal{E}^{1}_{\rho}(f_{2})\Big)d^{1}_{\rho}(f_{1},f_{2}).$
@@
We use the boundedness of $\mathcal{L}_{\delta}$ ((2.2), (2.4)) and of $H$ on $L^{2}(\mathbb{R})$ for estimates of $\mathcal{N}_{\delta}$ and those of $\mathcal{N}_{\infty}$, respectively. In the case of large $k$, we use appropriate interpolation inequalities to bound $\|\partial_{x}^{j}f\|_{L^{2}}$ and $\|\partial_{x}^{j}f\|_{L^{\infty}}$ for $1\leq j\leq k-1$ in terms of $\|\partial_{x}f\|_{L^{2}}$ and $\|\partial_{x}^{k}f\|_{L^{2}}$. A detailed verification is left to the reader. ∎
@@
We do not have a similar estimate for the term $\operatorname{Re}\big(f(iH)(f\partial_{x}\overline{g}),g\big)_{L^{2}}$, which suggests that the standard energy method is not sufficient to control this type of interactions. We will use the modified energy argument to handle this case.
@@
The first two terms are easily treated by integration by parts (one can estimate the second term without taking the real part). For the term involving $H$, an integration by parts yields
@@
Observe that the first term on the right-hand side vanishes by the positivity of $H\partial_{x}$, then we obtain the desired bound with no derivative on $g$. ∎

### v2

@@ 2.1. Notation
For $T>0$, we use the convention $L_{T}^{\infty}$ to denote the space $L^{\infty}(0,T)$. Furthermore, for a Banach space $B$, we use $L^{\infty}_{T}B$ to denote $L^{\infty}(0,T;B)$.
@@
We have already defined the space $\mathcal{Z}^{k}$, the distance function $d^{k}(\cdot,\cdot)$ and the functional $\mathcal{E}^{k}(\cdot)$ on it, in (1.2), (1.3) and (1.4), respectively. We also define the smooth subset $\mathcal{Z}^{\infty}:=\bigcap_{k\geq 1}\mathcal{Z}^{k}$.
@@ 2.2. Properties of $T_{\delta}$
Having an unbounded symbol (1.1), the operator $T_{\delta}$ itself is not bounded on $L^{p}(\mathbb{R})$ or $H^{s}(\mathbb{R})$. However, since the Fourier symbol $r_{\delta}(\xi)=(\coth(\delta\xi)-\mathrm{sgn}(\xi))\xi$ of $\mathcal{L}_{\delta}$ satisfies $|r_{\delta}(\xi)|\lesssim\delta^{-1}$, by Plancherel’s theorem we see that $\mathcal{L}_{\delta}$ is bounded on $H^{s}(\mathbb{R})$ for any $s\in\mathbb{R}$ and satisfies
@@
For $f\in\mathcal{Z}^{k}$, we notice that $\partial_{x}|f|^{2}=\partial_{x}(|f|^{2}-\rho_{f}^{2})$ gives
(2.4) $\|T_{\delta}\partial_{x}|f|^{2}\|_{L^{2}}$ $\lesssim\delta^{-1}\big\||f|^{2}-\rho_{f}^{2}\big\|_{L^{2}}+\|f\|_{L^{\infty}}\|\partial_{x}f\|_{L^{2}},$
$\|\mathcal{L}_{\delta}|f|^{2}\|_{L^{2}}$ $\lesssim\delta^{-1}\big\||f|^{2}-\rho_{f}^{2}\big\|_{L^{2}}.$
@@
2.3. Properties of the space $\mathcal{Z}^{k}$ and control of Littlewood-Paley operators
We begin this subsection with a basic estimate in $\mathcal{Z}^{k}$, which will be used in Section 4:
(2.6) $\rho\leq\|f\|_{L^{\infty}}\lesssim\rho+\big\||f|^{2}-\rho^{2}\big\|_{L^{2}}^{\frac{1}{3}}\|\partial_{x}f\|_{L^{2}}^{\frac{1}{3}}\qquad(f\in\mathcal{Z}^{1}_{\rho},~\rho\geq 0).$
The first inequality follows from $|f|^{2}-\rho^{2}\in L^{2}(\mathbb{R})$. The second inequality is obtained by Gagliardo-Nirenberg and Young’s inequalities as follows:
@@ 2.3. Properties of the space $\mathcal{Z}^{k}$ and control of Littlewood-Paley operators
Next, from Sobolev’s embedding $H^{1}(\mathbb{R})\hookrightarrow L^{\infty}(\mathbb{R})$ we see that for any integer $k\geq 1$, $\rho\geq 0$, $f\in\mathcal{Z}^{k}_{\rho}$ and $g\in H^{k}(\mathbb{R})$, it holds $f+g\in\mathcal{Z}^{k}_{\rho}$ and
(2.7) $\mathcal{E}^{k}(f+g)\lesssim\mathcal{E}^{k}(f)+\|g\|_{H^{k}}^{2}.$
In a similar manner (but using the Gagliardo-Nirenberg inequality instead of Sobolev’s embedding), for $f_{1},f_{2}\in\mathcal{Z}^{1}$ and $g_{1},g_{2}\in H^{1}(\mathbb{R})$, we can show that
(2.8) $d^{1}(f_{1}+g_{1},f_{2}+g_{2})$ $\lesssim\big(1+\|g_{1}\|_{L^{2}}\big)d^{1}(f_{1},f_{2})$
@@
Below, we collect some estimates for frequency localized functions. They are easily proven by using Bernstein’s inequality. For any $\rho\geq 0$, $\ell\geq 0$ and $k,j\geq 1$, we can show that $P_{\leq\ell}:\mathcal{Z}^{k}_{\rho}\to\mathcal{Z}^{k+j}_{\rho}$ with the estimates
(2.9) $\begin{aligned} \mathcal{E}^{k}(P_{\leq\ell}f)&\leq\mathcal{E}^{k}(f)+\big\||P_{\leq\ell}f|^{2}-|f|^{2}\big\|_{L^{2}}\leq\mathcal{E}^{k}(f)+2\|f\|_{L^{\infty}}\|P_{\geq\ell+1}f\|_{L^{2}}\\ &\leq\mathcal{E}^{k}(f)+2^{1-\ell}\|f\|_{L^{\infty}}\|\partial_{x}f\|_{L^{2}}\leq\mathcal{E}^{k}(f)+2^{1-\ell}\mathcal{E}^{1}(f)\leq 3\mathcal{E}^{k}(f),\end{aligned}$
$\mathcal{E}^{k+j}(P_{\leq\ell}f)\lesssim 2^{2j\ell}\mathcal{E}^{k}(f)+2^{-\ell}\mathcal{E}^{1}(f)\lesssim 2^{2j\ell}\mathcal{E}^{k}(f).$
We also see that $P_{\leq\ell}f\to f$ in $\mathcal{Z}^{k}$ as $\ell\to\infty$, thanks to the estimate
(2.10) $d^{k}(P_{\leq\ell}f,f)$ $\leq\|P_{\geq\ell+1}\partial_{x}f\|_{H^{k-1}}+\|P_{\geq\ell+1}f\|_{L^{\infty}}+\big\||P_{\leq\ell}f|^{2}-|f|^{2}\big\|_{L^{2}}$
@@
Note that if $f\in C([0,T];\mathcal{Z}^{k})$, then the above convergence is uniform in time on $[0,T]$. Using the identity
(2.11) $|Af|^{2}-|Ag|^{2}=|f|^{2}-|g|^{2}$ $-(1-A)(f-g)\cdot\overline{Af}-(1-A)g\cdot\overline{A(f-g)}$
$-(f-g)\cdot\overline{(1-A)f}-g\cdot\overline{(1-A)(f-g)}$
with $A=P_{\leq\ell}$, we have
(2.12) $\begin{aligned} d^{k}(P_{\leq\ell}f,P_{\leq\ell}g)&\leq d^{k}(f,g)\begin{aligned} &+\|P_{\geq\ell+1}(f-g)\|_{L^{2}}\|P_{\leq\ell}f\|_{L^{\infty}}+\|P_{\geq\ell+1}g\|_{L^{2}}\|P_{\leq\ell}(f-g)\|_{L^{\infty}}\\ &+\|f-g\|_{L^{\infty}}\|P_{\geq\ell+1}f\|_{L^{2}}+\|g\|_{L^{\infty}}\|P_{\geq\ell+1}(f-g)\|_{L^{2}}\end{aligned}\\ &\leq d^{k}(f,g)+2^{1-\ell}\Big(\big[\mathcal{E}^{1}(f)\big]^{\frac{1}{2}}+\big[\mathcal{E}^{1}(g)\big]^{\frac{1}{2}}\Big)d^{1}(f,g).\end{aligned}$
Finally, for different $\ell,m\in\mathbb{Z}_{\geq 0}$ and $f,g\in\mathcal{Z}^{2}$, we use the triangle inequality and (2.10) with $k=1$ to see that
(2.13) $d^{1}(P_{\leq\ell}f,P_{\leq m}g)\lesssim d^{1}(f,g)+2^{-\ell}\Big(\big[\mathcal{E}^{2}(f)\big]^{\frac{1}{2}}+\mathcal{E}^{2}(f)\Big)+2^{-m}\Big(\big[\mathcal{E}^{2}(g)\big]^{\frac{1}{2}}+\mathcal{E}^{2}(g)\Big).$
@@ 2.4. Estimates for the free Schrödinger operator
The action of $U(t)$ in the Zhidkov-type spaces has been well-studied; see, e.g., [18, 11].
@@
Using Gagliardo-Nirenberg inequality and (2.14), we see that
@@
(2.17) $U(t)\phi\in C(\mathbb{R};\mathcal{Z}^{k}_{\rho}),\quad\mathcal{E}^{k}(U(t)\phi)\lesssim(1+|t|^{\frac{3}{4}})\mathcal{E}^{k}(\phi)\qquad(\rho\geq 0,~\phi\in\mathcal{Z}^{k}_{\rho}).$
@@
(2.18) $d^{k}(U(t)\phi,U(t)\psi)\lesssim(1+|t|^{\frac{3}{4}})\Big(1+\mathcal{E}^{1}(\phi)+\mathcal{E}^{1}(\psi)\Big)^{\frac{1}{2}}d^{k}(\phi,\psi)$
for any $k\geq 1$ and $\phi,\psi\in\mathcal{Z}^{k}$. To prove (2.18), it is clearly sufficient to estimate the non-Sobolev part of $d^{1}$. From (2.15), we have
$\|U(t)\phi-U(t)\psi\|_{L^{\infty}}\lesssim\|\phi-\psi\|_{L^{\infty}}+|t|^{\frac{1}{4}}\|\partial_{x}(\phi-\psi)\|_{L^{2}}\lesssim(1+|t|^{\frac{1}{4}})d^{1}(\phi,\psi).$
Using the decomposition (2.11) with $A=U(t)$ and (2.14)–(2.15), we have
$\big\|(|U(t)\phi|^{2}-\rho_{\phi}^{2})-(|U(t)\psi|^{2}-\rho_{\psi}^{2})\big\|_{L^{2}}$
$\lesssim\big\|(|\phi|^{2}-\rho_{\phi}^{2})-(|\psi|^{2}-\rho_{\psi}^{2})\big\|_{L^{2}}\begin{aligned} &+\big\|U(t)(\phi-\psi)-(\phi-\psi)\big\|_{L^{2}}\|U(t)\phi\|_{L^{\infty}}+\|U(t)\psi-\psi\|_{L^{2}}\|U(t)(\phi-\psi)\|_{L^{\infty}}\\ &+\|\phi-\psi\|_{L^{\infty}}\|U(t)\phi-\phi\|_{L^{2}}+\|\psi\|_{L^{\infty}}\big\|U(t)(\phi-\psi)-(\phi-\psi)\big\|_{L^{2}}\end{aligned}$
$\lesssim(1+|t|^{\frac{3}{4}})\Big(1+\mathcal{E}^{1}(\phi)+\mathcal{E}^{1}(\psi)\Big)^{\frac{1}{2}}d^{1}(\phi,\psi),$
@@ 1. $\mathrm{(i)}$
    For $f\in\mathcal{Z}^{1}$,
    $\|\mathcal{N}_{\delta}(f)\|_{L^{2}}\lesssim\delta^{-1}\big[\mathcal{E}^{1}(f)\big]^{\frac{3}{2}}.$
    For any integer $k\geq 1$ and $f\in\mathcal{Z}^{k}$,
    $\|\partial_{x}^{k}\mathcal{N}_{\delta}(f)\|_{L^{2}}\lesssim_{k}\delta^{-1}\mathcal{E}^{1}(f)\|\partial_{x}f\|_{H^{k-1}}.$
@@ 2. $\mathrm{(ii)}$
    For any integer $k\geq 0$ and $f\in\mathcal{Z}^{k+1}_{*}$,
    $\|\partial_{x}^{k}\mathcal{N}_{\infty}(f)\|_{L^{2}}\lesssim_{k}\mathcal{E}^{1}(f)\|\partial_{x}f\|_{H^{k}}.$
@@ 3. $\mathrm{(iii)}$
    For $f_{1},f_{2}\in\mathcal{Z}^{1}$,
    $\delta\;\|\mathcal{N}_{\delta}(f_{1})-\mathcal{N}_{\delta}(f_{2})\|_{H^{1}}+\|\mathcal{N}_{\infty}(f_{1})-\mathcal{N}_{\infty}(f_{2})\|_{L^{2}}\lesssim\Big(1+\mathcal{E}^{1}(f_{1})+\mathcal{E}^{1}(f_{2})\Big)d^{1}(f_{1},f_{2}).$
@@
We use the boundedness of $\mathcal{L}_{\delta}$ ((2.2), (2.4)) and $H$ on $L^{2}(\mathbb{R})$ for estimates of $\mathcal{N}_{\delta}$ and $\mathcal{N}_{\infty}$, respectively. In the case $k\geq 2$, we use appropriate interpolation inequalities to bound $\|\partial_{x}^{j}f\|_{L^{2}}$ and $\|\partial_{x}^{j}f\|_{L^{\infty}}$ for $1\leq j\leq k-1$ in terms of $\|\partial_{x}f\|_{L^{2}}$ and $\|\partial_{x}^{k}f\|_{L^{2}}$. A detailed verification is left to the reader. ∎
@@
The right-hand side does not include $\partial_{x}g$. We do not have a similar estimate for the term $\operatorname{Re}\big(f(iH)(f\partial_{x}\overline{g}),g\big)_{L^{2}}$, which suggests that the standard energy method is not sufficient to control this type of interactions. We will use the modified energy method to handle this case.
@@
The first two terms are easily treated by integration by parts (one can estimate the second term without taking the real part). For the term involving $H$, we rewrite it as
@@
Observe that the first term on the right-hand side vanishes by the positivity of $H\partial_{x}$, then we obtain the desired bound. ∎